OpenStudy (anonymous):

A bag contains 10 red marbles, 15 yellow marbles and 5 green marbles. What is the probability of drawing a red or green marble from the bag?

OpenStudy (anonymous):

Probability of red*Probability of green

OpenStudy (anonymous):

They're mutually exclusive events, so you would add the probabilities.\[ \Pr(X=\text{red}) + \Pr(X=\text{green}) \]

OpenStudy (anonymous):

Since you draw only a single marble, it's not possible to draw both a red and green marble at the same time. This is what makes the mutually exclusive.

OpenStudy (anonymous):

In either case, there are \(10+15+5=30\) total marbles.

OpenStudy (anonymous):

\[ \Pr(X=\text{red}) = \frac{\text{number of red marbles}}{\text{total number of marbles}} \]

OpenStudy (anonymous):

so it would be 10/5

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Since \(5\) is not the total number of marbles.

OpenStudy (anonymous):

ohh so it would be 10/30

OpenStudy (anonymous):

Well, that is correct for pulling 1 red marble, so you're partially done.

OpenStudy (anonymous):

and for the pulling of a green marble it would be 5/30, correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok thanks soooo much

OpenStudy (anonymous):

What is the final answer?

OpenStudy (anonymous):

the answer choices were, A. 1/2, B. 1/4, C. 1/6, & D. 1/8

OpenStudy (anonymous):

I choose C

OpenStudy (anonymous):

becuase 5/30 simplified is 1/6

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

No, you need to add them. That gives you: \[ \frac{10}{30} + \frac{5}{30} = \frac{15}{30} = \frac 12 \]